# Backwards Data-Flow Analysis using Prophecy Variable in the BuildIt System

### v1

@@ 1. Introduction
GraphIt (g2) / GraphIt (BuildIt)(buildsl) 23,783 2,021 3%, 33%, -8%
COLA (shim) / SHIM (BuildIt) (shim) 7,500 5,700 1.4%, 32%, -30%
PCRE2(pcre2) / BREze (BuildIt) (tamara:meng-thesis:2023) 133,995 1,564 800%, 9100%, -19%
RE2(re2) / BREze (BuildIt) (tamara:meng-thesis:2023) 26,587 1,564 37%, 600%, -129%
Hyperscan(hyperscan) / BREze (BuildIt) (tamara:meng-thesis:2023) 187,033 1,564 -127%, 2000%, -1700%
@@
Component BuildIt(buildsl) GraphIt(g2)
@@ 1.1. BuildIt
In response to this situation, we developed BuildIt (buildsl; ajay-thesis), a system that leverages modern programming language features such as overloading to implement DSLs. Operating completely within a modern general purpose language, BuildIt implements compilation and optimization techniques that operate across the entire domain specific multioperation computation. BuildIt computations execute two stages. The first stage performs a combined partial evaluation, analysis, and compilation that generates optimized code . The second stage executes this optimized code (with the second stage execution typically comprising the overwhelming percentage of the overall computation). Because BuildIt is implemented as a collection of data structures implemented completely within a host general purpose language, it eliminates the need for language implementation components such as parsers and full blown intermediate representations. The result is a dramatic reduction in language implementation overhead. Table 2 quantifies this reduction by counting the number of lines of code in corresponding BuildIt and standard DSL implementations. Table 2 breaks down the lines of code for the GraphIt implementation components, highlighting the components that the lightweight BuildIt approach eliminates. These numbers show that BuildIt implementations can require up to 100 times fewer lines of code than standard implementations of the same DSL, a reduction substantial enough to transform the viability of many DSLs. And the performance results in Table 2 show that, in comparison with standard DSL implementations, BuildIt implementations typically exhibit comparable or better performance (buildsl; ajay-thesis).
@@ 1.2. Program Analysis in BuildIt
To apply the optimizations required to obtain acceptable performance, BuildIt must perform a range of program analyses. These take place during the BuildIt first phase and typically target the entire multioperation computation. For analyses that propagate information forward with the flow of control, BuildIt uses history variables (implemented as data structure fields in the underlying implementation language) to carry the information through the computation (OwickiG76; AbadiL91; rinard2020dataflow). These variables are assigned as the BuildIt first stage executes and carry the information through the first stage execution to be later accessed (again during the first stage) to optimize the generated second stage code.
@@ 1.3. Prophecy Variables, BuildIt, and Lightweight Program Analysis
Some optimizations, however, require information about the future program execution (history variables record information about the past, not future, execution). One standard way to obtain this information is to build an internal intermediate program representation, then use a backwards analysis to propagate information against the flow of control to reach the point where it is applied to optimize the generated code (DragonBook; appel2004modern; muchnick1997advanced; cooper2011engineering; kennedy2001optimizing). A drawback of attempting to apply this approach in the BuildIt context is the need to construct an appropriate program representation to enable the backwards analysis, an endeavor that would move the BuildIt implementation closer to standard DSLimplementations, increase the BuildIt implementation effort, and erode the implementation effort advantage that BuildIt currently exhibits over standard implementations of DSLs.
We instead present a novel use of prophecy variables to obtain information about the future program execution. Prophecy variables were initially developed to prove forward simulation relations in state machine refinement proofs that involve reasoning about future program behavior (AbadiL91). The basic idea behind prophecy variables in the BuildIt context is to use prophecy variable predictions to predict information about the future execution, then, when the information becomes available later in the execution, use prophecy variable preconditions to check that the prophecy variable correctly predicts the information.
In the refinement proof context, prophecy variable predictions are typically nondeterministic, with angelic nondeterminism enabling the refinement proof to go through if any of the nondeterministic predictions was correct (AbadiL91). In the BuildIt context, we instead use repeated program executions to discover correct prophecy variable predictions. When the BuildIt implementation encounters a violated prophecy variable precondition (this occurs when there is an incorrect prophecy variable prediction), it updates the prophecy variable to satisfy the precondition, propagates the update to the prophecy variable prediction point, terminates the BuildIt (first phase) execution, then reexecutes with the new prophecy variable prediction, repeating this process until all prophecy variable predictions are correct.
@@ 2. Example
Figure 1 presents example BuildIt code for tensor computations using the Einstein summation notation (einstein_foundation_1916) (einsum notation) written in C++ (einsum-buildit). The example uses prophecy variables to predict tensors that may be accessed by a computation performed on a GPU. This information enables BuildIt to generate code that moves the accessed tensors to GPU memory before executing the GPU computation. The example uses several BuildIt types that guide the BuildIt staged execution. static_var and prophecy_var variables are first stage variables that exist only during the first stage (with the values of updated prophecy_var variables persisting across first stage reexecutions triggered by violated prophecy variable preconditions). dyn_var variables are second stage variables whose values are computed only during the second stage. Expressions and control flow that depend only on the values of static_var and prophecy_var variables are fully computed during the first stage. The first stage produces (optimized) second stage code that computes second stage expressions and control flow that depends on the values of dyn_var variables. Second stage expressions may contain first stage variables. BuildIt replaces these first stage variables in the generated second stage code with the corresponding first stage values that exist during first stage code generation, effectively specializing the generated second stage code to the first stage values.
@@ 3.1. Overview
We formalize prophecy variables and the prophecy variable analysis using a core programming language inspired by Glynn Winskell’s imperative language IMP (Winskel93). Notable differences include the introduction of labels for all commands and the use of variables $V$ instead of locations ($\mathbf{Loc}$(rinard2020dataflow). Specifically, the formalization works with programs $P$ that contain labeled commands of the form $l:c\in P$, where $l,g\in L$, $c\in C$. $\textbf{labels}(P)=\{l.l:c\in P\}$ is the set of labels in $P$. Labels are unique — no two labeled commands in $P$ have the same label $l$. An executing program operates on states $\sigma\in\Sigma$ that maps variables $v,w\in V$ to integer values $n,m\in N$. The standard operational semantics is modeled by a program execution transition relation $\left\langle l,\sigma\right\rangle\rightarrow\left\langle l^{\prime},\sigma^{\prime}\right\rangle$. Execution starts at $\left\langle l_{0},\sigma_{0}\right\rangle$, where $l_{0}=\mathbf{first}(P)$ is the label of the first command to execute and $\sigma_{0}$ is the initial state. Execution terminates if it encounters an $l:\mathbf{halt}$ command.
@@
We note that if Preservation (Definition 3.1) and Progress (Definition 3.2) both hold, then the relation $\sim$ defined by $\left\langle l,\sigma\right\rangle\sim\left\langle l,\sigma,\beta_{\bullet l}\right\rangle$ is a bisimulation relation (Milner89; Benthem96) between standard and extended program configurations (rinard2020dataflow).
@@ 3.5. Prophecy Variable Analysis via Repeated Program Execution
It is possible to use a backwards dataflow analyses to obtain prophecy variable values that correctly predict future program executions (rinard2020dataflow). These analyses, however, work with an explicit program representation to propagate information backwards against the flow of control. They are therefore incompatible with the BuildIt staged compilation approach, which executes within a standard forward program execution environment to eliminate language implementation components such as intermediate representations. We therefore use an alternate approach that uses repeated program reexecutions to compute prophecy variable values.
@@ 4.1. BuildIt Overview
BuildIt is a lightweight DSL implementation system (ajay-thesis). BuildIt programs execute two stages. The first stage executes as standard C++ code that generates optimized C, C++, and/or CUDA code that executes in the second stage. In a typical BuildIt program the vast majority of the computation is performed by the second stage.
@@ 4.2. Prophecy Variables in BuildIt
BuildIt also uses program reexecution to explore and generate control flow that involves dyn_var<T> expressions (buildsl; ajay-thesis). We repurpose this program reexecution capability to implement the prophecy variable analysis formalized in Section 3.
@@ 4.3. Tensor Data Movement for GPU Kernels
This benchmark involves a BuildIt DSL that can evaluate arbitrary multidimensional tensor operations specified as einsum expressions (einsum-buildit). In this DSL the the dimensions and sizes of each tensor are first stage values. As a BuildIt DSL, the DSL is implemented as a C++ data structure library. The first stage executes with the standard C++ execution environment to generate second stage tensor code, including the generation of CUDA code for GPU kernels. The library introduces the key type Tensor<E>, where the first stage template parameter E is the element type of the tensor. The first stage tensor dimensions are stored internally as static_var<std::vector<int>> members. The second stage tensor values are stored in members of type dyn_var<E*>. Tensor operations such as [], +, *, =, += and *= are overloaded to implement the einsum operations.
@@ 9 C[i][j] += A[i][k] * B[k][j];
We present performance results for two computations: 1) matrix-matrix multiply (six tensors, three of which are ever resident on the GPU) and 2) matrix-vector multiply (two matrices and four vectors, with one matrix and one vector ever resident on the GPU). We consider two alternative GPU tensor movement implementations: 1) moving all tensors to the GPU before executing the GPU kernel (we acknowledge that this strategy becomes infeasible for applications that manipulate large numbers of tensors) and 2) using a runtime data movement strategy ( CUDA unified memory, which uses a page fault based mechanism to move data between CPU and GPU (Harris2017UnifiedMemory)). Table 8 presents the minimum execution time in microseconds for GPU matrix-matrix multiplication running with the three strategies and four matrix sizes across ten runs. Table 8 presents the minimum execution time in microseconds for GPU matrix-vector multiplication running with the three strategies and four matrix sizes across ten runs. The hardware platform for these experiments is an NVIDIA DGX-1 with an Intel(R) Xeon(R) CPU E5-2698 v4 @ 2.20GHz and a Tesla V100-SXM2-32GB GPU. BuildIt with prophecy variables always exhibits the best performance, with the relative performance gap narrowing as the matrix size increases. We attribute this phenomenon to the fact that, because matrix multiply performs $O(n^{3})$ computation and $O(n^{2})$ communication, at larger matrix sizes the computation time increasingly dominates so that the communication mechanism plays a smaller role in the overall performance.
@@
Table 9 presents the number of first stage program executions required to 1) fully explore all possible dynamic control flow paths (this is a baseline BuildIt technique for ensuring correct second stage code generation (buildit; ajay-thesis)) and 2) obtain correct prophecy variable values. It also presents the first stage execution times. Note that the use of prophecy variables can increase the number of dynamic control flow paths that BuildIt must explore to generate correct second stage code, which in turn causes an increase in the number of dynamic control flow exploration executions with prophecy variables and a corresponding increase in first stage execution times.
@@
Figure 9. Number of first stage program executions and first stage execution times for Matrix Matrix Multiply and Matrix Vector Multiply. We report both dynamic control flow exploration executions (ajay-thesis) and executions triggered by prophecy variable precondition violations.
@@ 4.4. Neural Network Operation Fusion
    Convolution-ReLU Fusion Fusion Disabled Enabled Prophecy Variable Runs - 4 Dynamic Control Flow Runs 33 72 First Stage Execution Time 9,872 us 14,701 us Figure 12. Number of first stage program executions and first stage execution times. We report both dynamic control flow exploration executions (ajay-thesis) and executions triggered by prophecy variable precondition violations.
@@ 5.1. Prophecy and History Variables
Many of the concepts that appear in simulation relation proofs for state machines also appear in the program verification, dataflow analysis, and abstract interpretation literature. For example, history variables were first introduced in the program verification literature (OwickiG76), abstraction functions, originally introduced in the program verification literature (Hoare72), can be seen as a form of refinement mappings, and program analyses can be seen as establishing a simulation relation between an abstract interpretation of the program (which plays the role of the specification) and concrete executions of the program (which play the role of the implementation) (CousotC77; CousotC92). It is also known that, in this context, backward or reverse simulation relations can be used to establish the correspondence between backward analyses (which extract information about the future execution) and program executions (CousotC92; SchmidtS98).
@@
A recent (unpublished in any peer reviewed venue) manuscript introduces prophecy variables to enable forward reasoning about program analysis properties that involve the future execution of the program (rinard2020dataflow). To the best of our knowledge, this manuscript is the first to introduce prophecy variables for program analysis (here we contrast with the use of prophecy variables for program verification (JungLPRTDJ20; Vafeiadis08; ZhangFFSL12) as well as the traditional use of prophecy variables for proving forward simulation relations between state machines (AbadiL91)). The manuscript uses prophecy variables to specify two program analysis problems, live variables and very busy expressions, that require information about the future execution of the program and uses a backwards dataflow analysis to obtain correct prophecy variable predictions.
@@
While we build on the formalism presented in this manuscript, our research differs in that 1) we use repeated forward program executions, not backwards dataflow analysis, to find correct prophecy variable predictions, 2) we implement prophecy variables in the BuildIt staged DSL implementation system ( (rinard2020dataflow) presents no implementation), 3) we show how prophecy variables enable the BuildIt first stage, which uses only standard forward program execution, to obtain and exploit information about the future execution of the program, and 4) we report results from BuildIt prophecy variable implementations. The overall result is the extension of the scope of the lightweight BuildIt implementation approach to include optimizations (such as those presented in Section 4) that require information about the future execution of the program.
@@
Cobalt enables compiler developers to specify a range of dataflow optimizations (such as constant propagation and partial dead assignment elimination) (LernerMC03). Each optimization is specified by a transformation pattern whose guard specifies a condition over sequences of actions in paths in the program representation that must hold for the transformation to be legal. Cobalt has separate constructs for specifying forward and backward optimizations — forward guards reason about forward properties, backward guards reason about backward properties.
@@
Rhodium was developed to eliminate the complexity of using temporal logic formulas and implicit dataflow facts to reason about paths in control-flow graphs (LernerMRC05). Rhodium uses de facto abstraction functions (expressed as predicates over concrete program states) to establish the connection between concrete program states and dataflow facts and state extensions (a form of instrumented semantics) to support analyses that extract information about the past execution of the program not present in standard concrete program states. Like Cobalt, Rhodium has separate support for forward and backward analyses; subsequent work on automatically inferring correct propagation rules supports only forward rules (ScherpelzLC07).
@@
Simulation relations, and techniques for proving that simulation relations exist, have been extensively explored in the context of establishing simulation relations between state machines (LynchV95; LynchDistributedAlgorithms). The developed theory includes a range of proof techniques and mechanisms, including forward and backward proof techniques with refinement mappings, abstraction functions, and abstraction relations. Prophecy variables were initially developed for the purpose of proving that implementations satisfy specifications via refinement mappings with forward simulations, specifically in the case when the specification makes a choice before the implementation (AbadiL91). The addition of prophecy variables to the framework of refinement mappings with history variables and forward simulation proofs enabled a completeness result for the ability to prove trace inclusions of implementations within specifications (AbadiL91). It is, of course, known that backward simulation is an alternative to forward simulation with prophecy variables (LynchV95). In general, there are a number of alternatives when choosing a formal framework for proving simulation properties, with the appropriate framework depending on pragmatic issues such as the convenience and conceptual difficulty of working with the concepts in the framework. In general, approaches that reason forward in time seem to be more attractive and intuitive than approaches that reason backward against time, as can be seen, for example, in pedagogical presentations of dataflow analyses, which invariably present forward analyses first, then backward analyses second as a kind of dual of forward analyses (appel2004modern; muchnick1997advanced; DragonBook; cooper2011engineering; kennedy2001optimizing; Aldrich2019Correctness; Aldrich2019Examples).
Prophecy variables have recently been applied for program verification in a Hoare program logic based on separation logic (JungLPRTDJ20). They have also been used as a mechanism to support reasoning about transformations between different program representations (LalR08; CookK11). Our purpose is different, specifically to use prophecy variables to enable forward reasoning and eliminate the need for backwards analyses and explicit intermediate representations in the context of a lightweight DSL implementation system that uses information about the future execution of the program to optimize the generated code.

### v2

# Backwards Data-Flow Analysis using Prophecy Variables in the BuildIt System

@@ 1. Introduction
GraphIt (Brahmakshatriya et al., 2021b) / GraphIt (BuildIt)(Brahmakshatriya and Amarasinghe, 2022) 23,783 2,021 3%, 33%, -8%
COLA (Ray, 2023) / SHIM (BuildIt) (Ray, 2023) 7,500 5,700 1.4%, 32%, -30%
PCRE2(Project, 2025) / BREze (BuildIt) (Mitrovska, 2023) 133,995 1,564 800%, 9100%, -19%
RE2(Cox, 2009) / BREze (BuildIt) (Mitrovska, 2023) 26,587 1,564 37%, 600%, -129%
Hyperscan(Wang et al., 2019) / BREze (BuildIt) (Mitrovska, 2023) 187,033 1,564 -127%, 2000%, -1700%
@@
Component BuildIt(Brahmakshatriya and Amarasinghe, 2022) GraphIt(Brahmakshatriya et al., 2021b)
@@ 1.1. BuildIt
In response to this situation, we developed BuildIt (Brahmakshatriya and Amarasinghe, 2022; Brahmakshatriya, 2025), a system that leverages modern programming language features such as overloading to implement DSLs. Operating completely within a modern general purpose language, BuildIt implements compilation and optimization techniques that operate across the entire domain specific multioperation computation. BuildIt computations execute two stages. The first stage performs a combined partial evaluation, analysis, and compilation that generates optimized code . The second stage executes this optimized code (with the second stage execution typically comprising the overwhelming percentage of the overall computation). Because BuildIt is implemented as a collection of data structures implemented completely within a host general purpose language, it eliminates the need for language implementation components such as parsers and full blown intermediate representations. The result is a dramatic reduction in language implementation overhead. Table 2 quantifies this reduction by counting the number of lines of code in corresponding BuildIt and standard DSL implementations. Table 2 breaks down the lines of code for the GraphIt implementation components, highlighting the components that the lightweight BuildIt approach eliminates. These numbers show that BuildIt implementations can require up to 100 times fewer lines of code than standard implementations of the same DSL, a reduction substantial enough to transform the viability of many DSLs. And the performance results in Table 2 show that, in comparison with standard DSL implementations, BuildIt implementations typically exhibit comparable or better performance (Brahmakshatriya and Amarasinghe, 2022; Brahmakshatriya, 2025).
@@ 1.2. Program Analysis in BuildIt
To apply the optimizations required to obtain acceptable performance, BuildIt must perform a range of program analyses. These take place during the BuildIt first phase and typically target the entire multioperation computation. For analyses that propagate information forward with the flow of control, BuildIt uses history variables (implemented as data structure fields in the underlying implementation language) to carry the information through the computation (Owicki and Gries, 1976; Abadi and Lamport, 1991; Rinard and Gadient, 2020). These variables are assigned as the BuildIt first stage executes and carry the information through the first stage execution to be later accessed (again during the first stage) to optimize the generated second stage code.
@@ 1.3. Prophecy Variables, BuildIt, and Lightweight Program Analysis
Some optimizations, however, require information about the future program execution (history variables record information about the past, not future, execution). One standard way to obtain this information is to build an internal intermediate program representation, then use a backwards analysis to propagate information against the flow of control to reach the point where it is applied to optimize the generated code (Aho et al., 2006; Appel, 2004; Muchnick and others, 1997; Cooper and Torczon, 2011; Kennedy and Allen, 2001). A drawback of attempting to apply this approach in the BuildIt context is the need to construct an appropriate program representation to enable the backwards analysis, an endeavor that would move the BuildIt implementation closer to standard DSLimplementations, increase the BuildIt implementation effort, and erode the implementation effort advantage that BuildIt currently exhibits over standard implementations of DSLs.
We instead present a novel use of prophecy variables to obtain information about the future program execution. Prophecy variables were initially developed to prove forward simulation relations in state machine refinement proofs that involve reasoning about future program behavior (Abadi and Lamport, 1991). The basic idea behind prophecy variables in the BuildIt context is to use prophecy variable predictions to predict information about the future execution, then, when the information becomes available later in the execution, use prophecy variable preconditions to check that the prophecy variable correctly predicts the information.
In the refinement proof context, prophecy variable predictions are typically nondeterministic, with angelic nondeterminism enabling the refinement proof to go through if any of the nondeterministic predictions was correct (Abadi and Lamport, 1991). In the BuildIt context, we instead use repeated program executions to discover correct prophecy variable predictions. When the BuildIt implementation encounters a violated prophecy variable precondition (this occurs when there is an incorrect prophecy variable prediction), it updates the prophecy variable to satisfy the precondition, propagates the update to the prophecy variable prediction point, terminates the BuildIt (first phase) execution, then reexecutes with the new prophecy variable prediction, repeating this process until all prophecy variable predictions are correct.
@@ 2. Example
Figure 1 presents example BuildIt code for tensor computations using the Einstein summation notation (Einstein, 1916) (einsum notation) written in C++ (BuildIt-lang, 2022). The example uses prophecy variables to predict tensors that may be accessed by a computation performed on a GPU. This information enables BuildIt to generate code that moves the accessed tensors to GPU memory before executing the GPU computation. The example uses several BuildIt types that guide the BuildIt staged execution. static_var and prophecy_var variables are first stage variables that exist only during the first stage (with the values of updated prophecy_var variables persisting across first stage reexecutions triggered by violated prophecy variable preconditions). dyn_var variables are second stage variables whose values are computed only during the second stage. Expressions and control flow that depend only on the values of static_var and prophecy_var variables are fully computed during the first stage. The first stage produces (optimized) second stage code that computes second stage expressions and control flow that depends on the values of dyn_var variables. Second stage expressions may contain first stage variables. BuildIt replaces these first stage variables in the generated second stage code with the corresponding first stage values that exist during first stage code generation, effectively specializing the generated second stage code to the first stage values.
@@ 3.1. Overview
We formalize prophecy variables and the prophecy variable analysis using a core programming language inspired by Glynn Winskell’s imperative language IMP (Winskel, 1993). Notable differences include the introduction of labels for all commands and the use of variables $V$ instead of locations ($\mathbf{Loc}$(Rinard and Gadient, 2020). Specifically, the formalization works with programs $P$ that contain labeled commands of the form $l:c\in P$, where $l,g\in L$, $c\in C$. $\textbf{labels}(P)=\{l.l:c\in P\}$ is the set of labels in $P$. Labels are unique — no two labeled commands in $P$ have the same label $l$. An executing program operates on states $\sigma\in\Sigma$ that maps variables $v,w\in V$ to integer values $n,m\in N$. The standard operational semantics is modeled by a program execution transition relation $\left\langle l,\sigma\right\rangle\rightarrow\left\langle l^{\prime},\sigma^{\prime}\right\rangle$. Execution starts at $\left\langle l_{0},\sigma_{0}\right\rangle$, where $l_{0}=\mathbf{first}(P)$ is the label of the first command to execute and $\sigma_{0}$ is the initial state. Execution terminates if it encounters an $l:\mathbf{halt}$ command.
@@
We note that if Preservation (Definition 3.1) and Progress (Definition 3.2) both hold, then the relation $\sim$ defined by $\left\langle l,\sigma\right\rangle\sim\left\langle l,\sigma,\beta_{\bullet l}\right\rangle$ is a bisimulation relation (Milner, 1989; van Benthem, 1996) between standard and extended program configurations (Rinard and Gadient, 2020).
@@ 3.5. Prophecy Variable Analysis via Repeated Program Execution
It is possible to use a backwards dataflow analyses to obtain prophecy variable values that correctly predict future program executions (Rinard and Gadient, 2020). These analyses, however, work with an explicit program representation to propagate information backwards against the flow of control. They are therefore incompatible with the BuildIt staged compilation approach, which executes within a standard forward program execution environment to eliminate language implementation components such as intermediate representations. We therefore use an alternate approach that uses repeated program reexecutions to compute prophecy variable values.
@@ 4.1. BuildIt Overview
BuildIt is a lightweight DSL implementation system (Brahmakshatriya, 2025). BuildIt programs execute two stages. The first stage executes as standard C++ code that generates optimized C, C++, and/or CUDA code that executes in the second stage. In a typical BuildIt program the vast majority of the computation is performed by the second stage.
@@ 4.2. Prophecy Variables in BuildIt
BuildIt also uses program reexecution to explore and generate control flow that involves dyn_var<T> expressions (Brahmakshatriya and Amarasinghe, 2022; Brahmakshatriya, 2025). We repurpose this program reexecution capability to implement the prophecy variable analysis formalized in Section 3.
@@ 4.3. Tensor Data Movement for GPU Kernels
This benchmark involves a BuildIt DSL that can evaluate arbitrary multidimensional tensor operations specified as einsum expressions (BuildIt-lang, 2022). In this DSL the the dimensions and sizes of each tensor are first stage values. As a BuildIt DSL, the DSL is implemented as a C++ data structure library. The first stage executes with the standard C++ execution environment to generate second stage tensor code, including the generation of CUDA code for GPU kernels. The library introduces the key type Tensor<E>, where the first stage template parameter E is the element type of the tensor. The first stage tensor dimensions are stored internally as static_var<std::vector<int>> members. The second stage tensor values are stored in members of type dyn_var<E*>. Tensor operations such as [], +, *, =, += and *= are overloaded to implement the einsum operations.
@@ 9 C[i][j] += A[i][k] * B[k][j];
We present performance results for two computations: 1) matrix-matrix multiply (six tensors, three of which are ever resident on the GPU) and 2) matrix-vector multiply (two matrices and four vectors, with one matrix and one vector ever resident on the GPU). We consider two alternative GPU tensor movement implementations: 1) moving all tensors to the GPU before executing the GPU kernel (we acknowledge that this strategy becomes infeasible for applications that manipulate large numbers of tensors) and 2) using a runtime data movement strategy ( CUDA unified memory, which uses a page fault based mechanism to move data between CPU and GPU (Harris, 2017)). Table 8 presents the minimum execution time in microseconds for GPU matrix-matrix multiplication running with the three strategies and four matrix sizes across ten runs. Table 8 presents the minimum execution time in microseconds for GPU matrix-vector multiplication running with the three strategies and four matrix sizes across ten runs. The hardware platform for these experiments is an NVIDIA DGX-1 with an Intel(R) Xeon(R) CPU E5-2698 v4 @ 2.20GHz and a Tesla V100-SXM2-32GB GPU. BuildIt with prophecy variables always exhibits the best performance, with the relative performance gap narrowing as the matrix size increases. We attribute this phenomenon to the fact that, because matrix multiply performs $O(n^{3})$ computation and $O(n^{2})$ communication, at larger matrix sizes the computation time increasingly dominates so that the communication mechanism plays a smaller role in the overall performance.
@@
Table 9 presents the number of first stage program executions required to 1) fully explore all possible dynamic control flow paths (this is a baseline BuildIt technique for ensuring correct second stage code generation (Brahmakshatriya and Amarasinghe, 2021; Brahmakshatriya, 2025)) and 2) obtain correct prophecy variable values. It also presents the first stage execution times. Note that the use of prophecy variables can increase the number of dynamic control flow paths that BuildIt must explore to generate correct second stage code, which in turn causes an increase in the number of dynamic control flow exploration executions with prophecy variables and a corresponding increase in first stage execution times.
@@
Figure 9. Number of first stage program executions and first stage execution times for Matrix Matrix Multiply and Matrix Vector Multiply. We report both dynamic control flow exploration executions (Brahmakshatriya, 2025) and executions triggered by prophecy variable precondition violations.
@@ 4.4. Neural Network Operation Fusion
    Convolution-ReLU Fusion Fusion Disabled Enabled Prophecy Variable Runs - 4 Dynamic Control Flow Runs 33 72 First Stage Execution Time 9,872 us 14,701 us Figure 12. Number of first stage program executions and first stage execution times. We report both dynamic control flow exploration executions (Brahmakshatriya, 2025) and executions triggered by prophecy variable precondition violations.
@@ 5.1. Prophecy and History Variables
Many of the concepts that appear in simulation relation proofs for state machines also appear in the program verification, dataflow analysis, and abstract interpretation literature. For example, history variables were first introduced in the program verification literature (Owicki and Gries, 1976), abstraction functions, originally introduced in the program verification literature (Hoare, 1972), can be seen as a form of refinement mappings, and program analyses can be seen as establishing a simulation relation between an abstract interpretation of the program (which plays the role of the specification) and concrete executions of the program (which play the role of the implementation) (Cousot and Cousot, 1977, 1992). It is also known that, in this context, backward or reverse simulation relations can be used to establish the correspondence between backward analyses (which extract information about the future execution) and program executions (Cousot and Cousot, 1992; Schmidt and Steffen, 1998).
@@
A recent (unpublished in any peer reviewed venue) manuscript introduces prophecy variables to enable forward reasoning about program analysis properties that involve the future execution of the program (Rinard and Gadient, 2020). To the best of our knowledge, this manuscript is the first to introduce prophecy variables for program analysis (here we contrast with the use of prophecy variables for program verification (Jung et al., 2020; Vafeiadis, 2008; Zhang et al., 2012) as well as the traditional use of prophecy variables for proving forward simulation relations between state machines (Abadi and Lamport, 1991)). The manuscript uses prophecy variables to specify two program analysis problems, live variables and very busy expressions, that require information about the future execution of the program and uses a backwards dataflow analysis to obtain correct prophecy variable predictions.
@@
While we build on the formalism presented in this manuscript, our research differs in that 1) we use repeated forward program executions, not backwards dataflow analysis, to find correct prophecy variable predictions, 2) we implement prophecy variables in the BuildIt staged DSL implementation system ( (Rinard and Gadient, 2020) presents no implementation), 3) we show how prophecy variables enable the BuildIt first stage, which uses only standard forward program execution, to obtain and exploit information about the future execution of the program, and 4) we report results from BuildIt prophecy variable implementations. The overall result is the extension of the scope of the lightweight BuildIt implementation approach to include optimizations (such as those presented in Section 4) that require information about the future execution of the program.
@@
Cobalt enables compiler developers to specify a range of dataflow optimizations (such as constant propagation and partial dead assignment elimination) (Lerner et al., 2003). Each optimization is specified by a transformation pattern whose guard specifies a condition over sequences of actions in paths in the program representation that must hold for the transformation to be legal. Cobalt has separate constructs for specifying forward and backward optimizations — forward guards reason about forward properties, backward guards reason about backward properties.
@@
Rhodium was developed to eliminate the complexity of using temporal logic formulas and implicit dataflow facts to reason about paths in control-flow graphs (Lerner et al., 2005). Rhodium uses de facto abstraction functions (expressed as predicates over concrete program states) to establish the connection between concrete program states and dataflow facts and state extensions (a form of instrumented semantics) to support analyses that extract information about the past execution of the program not present in standard concrete program states. Like Cobalt, Rhodium has separate support for forward and backward analyses; subsequent work on automatically inferring correct propagation rules supports only forward rules (Scherpelz et al., 2007).
@@
Simulation relations, and techniques for proving that simulation relations exist, have been extensively explored in the context of establishing simulation relations between state machines (Lynch and Vaandrager, 1995; Lynch, 1996). The developed theory includes a range of proof techniques and mechanisms, including forward and backward proof techniques with refinement mappings, abstraction functions, and abstraction relations. Prophecy variables were initially developed for the purpose of proving that implementations satisfy specifications via refinement mappings with forward simulations, specifically in the case when the specification makes a choice before the implementation (Abadi and Lamport, 1991). The addition of prophecy variables to the framework of refinement mappings with history variables and forward simulation proofs enabled a completeness result for the ability to prove trace inclusions of implementations within specifications (Abadi and Lamport, 1991). It is, of course, known that backward simulation is an alternative to forward simulation with prophecy variables (Lynch and Vaandrager, 1995). In general, there are a number of alternatives when choosing a formal framework for proving simulation properties, with the appropriate framework depending on pragmatic issues such as the convenience and conceptual difficulty of working with the concepts in the framework. In general, approaches that reason forward in time seem to be more attractive and intuitive than approaches that reason backward against time, as can be seen, for example, in pedagogical presentations of dataflow analyses, which invariably present forward analyses first, then backward analyses second as a kind of dual of forward analyses (Appel, 2004; Muchnick and others, 1997; Aho et al., 2006; Cooper and Torczon, 2011; Kennedy and Allen, 2001; Aldrich, 2019b, a).
Prophecy variables have recently been applied for program verification in a Hoare program logic based on separation logic (Jung et al., 2020). They have also been used as a mechanism to support reasoning about transformations between different program representations (Lal and Reps, ; Cook and Koskinen, ). Our purpose is different, specifically to use prophecy variables to enable forward reasoning and eliminate the need for backwards analyses and explicit intermediate representations in the context of a lightweight DSL implementation system that uses information about the future execution of the program to optimize the generated code.